### v1

# Fine-grained Image-to-LiDAR Contrastive Distillation with Visual Foundation Models

Abstract
Contrastive image-to-LiDAR knowledge transfer, commonly used for learning 3D representations with synchronized images and point clouds, often faces a self-conflict dilemma. This issue arises as contrastive losses unintentionally dissociate features of unmatched points and pixels that share semantic labels, compromising the integrity of learned representations. To overcome this, we harness Visual Foundation Models (VFMs), which have revolutionized the acquisition of pixel-level semantics, to enhance 3D representation learning. Specifically, we utilize off-the-shelf VFMs to generate semantic labels for weakly-supervised pixel-to-point contrastive distillation. Additionally, we employ von Mises-Fisher distributions to structure the feature space, ensuring semantic embeddings within the same class remain consistent across varying inputs. Furthermore, we adapt sampling probabilities of points to address imbalances in spatial distribution and category frequency, promoting comprehensive and balanced learning. Extensive experiments demonstrate that our approach mitigates the challenges posed by traditional methods and consistently surpasses existing image-to-LiDAR contrastive distillation methods in downstream tasks. The source code is available at https://github.com/Eaphan/OLIVINE.
1 Introduction
LiDAR sensors deliver critical information in the 3D environment, crucial for applications such as autonomous driving [66; 63]. State-of-the-art neural networks have shown promising performance on point clouds processing, which rely on extensive annotated datasets [25; 65; 67]. However, the process of annotating point clouds is not only time-consuming but also costly, presenting significant challenges in terms of scalability and practicality [31]. Self-supervision offers a solution by leveraging vast quantities of unlabeled data to pre-train networks and subsequently fine-tuning them with a smaller, labeled dataset. This approach significantly reduces the reliance on extensive annotated data sets [9].
A prevalent method for learning 3D representations involves contrastive pixel-to-point knowledge transfer, using synchronized and calibrated images and point clouds. PPKT [34] enables a 3D network to derive extensive knowledge from a pre-trained 2D image backbone through a pixel-to-point contrastive loss. This entire pre-training process necessitates no annotations for either images or point clouds. Then SLidR [43] employs superpixels to cluster pixels and points from visually coherent regions, leading to a more meaningful contrastive task. Building on this, Seal [33] utilizes semantically rich superpixels generated by visual foundation models and introduces temporal consistency regularization across point segments at different times. Meanwhile, HVDistill [62] innovates by implementing cross-modality contrastive distillation that integrates both image-plane and bird-eye views.
Unfortunately, existing contrastive distillation methods are hindered by several critical limitations. Firstly, a "self-conflict" issue arises during pre-training, where (super)pixels that belong to the same category as an anchor (super)point but do not directly correspond are simply treated as negative samples (see Fig.1(a)). This approach neglects the inherent semantic connections within the same category, leading to conflicts in the learning process where beneficial relationships might be disregarded. This problem is magnified by the contrastive loss’s intrinsic hardness-aware characteristic, which results in the most substantial gradient influences derived from negative samples that are semantically the most similar [50; 36]. While ST-SLidR [36] has introduced a semantically tolerant loss to mitigate this issue, the absence of a robust high-level semantic understanding could not fundamentally change
the intrinsic hardness-aware nature of the contrastive loss. Secondly, conventional sampling methods for point-pixel pairs fail to consider significant category imbalances or variations in point densities relative to their distance from sensors [34]. For example, bicycles comprise only 1.47% of annotations in the nuScenes-lidarseg dataset, whereas drivable surfaces make up 37.66%. This oversight can result in a skewed representation of the environment, where dominant categories or densely populated areas are over-represented, impacting the model’s effectiveness and fairness.
In this study, we address the "self-conflict" issue by leveraging supervised contrastive distillation enhanced with weak semantic labels generated by VFMs (Visual Foundation Models). VFMs like
Figure 1: Illustration of (a) self-conflict that exists in conventional pixel-to-point contrastive distillation and (b) our weakly supervised contrastive distillation.
SAM (Segment Anything Model), trained on expansive datasets, has revolutionized computer vision by simplifying the acquisition of pixel-level semantics. These models are exceptionally adaptable, enabling the direct derivation of semantic labels through specified prompts without the need for retraining. As depicted in Fig.1(b), with these weak labels, we draw the embeddings of an anchor point and corresponding pixels of the same class closer together, while distancing the anchor from “negative” pixels of differing classes. Furthermore, given that representation of samples in the same class can vary significantly across different batches, we introduce semantic-guided consistency regularization to enhance 3D representation learning. This approach structures the feature space by modeling each class with a von Mises-Fisher distribution and making point features adhere closely to their respective distributions.
Considering the challenges posed by category imbalance and the non-uniform distribution of point clouds, we propose a density and category-aware sampling strategy. This method accounts for both the density of points and the frequency of their categories. By adjusting the sampling probabilities of different anchor points, we enhance the quality of learned 3D representations, particularly for points that fall into minority categories or are situated in areas of low density.
Extensive experiments reveal that our pre-training approach outperforms state-of-the-art 3D self-supervised methods on both the nuScenes and SemanticKITTI datasets. The primary contributions of this work are summarized as follows: 1) To tackle the challenge of “self-conflict”, we utilize off-the-shelf VFMs to generate semantic labels for weakly-supervised pixel-to-point contrastive distillation. 2) We introduce semantic-guided consistency regularization to cultivate a meaningful and structured feature space. 3) We develop an innovative sampling strategy for point-pixel pairs that considers both the category frequency and spatial density of points.
2 Related Work
3D Representation Learning. Recent advancements in 3D self-supervised learning have closely paralleled innovations in the image domain, extending these methods to diverse 3D contexts such as object-level point clouds [42; 51; 21], indoor scenes [52; 11; 69; 57; 8; 27; 58; 49], and outdoor settings [29; 3; 60; 37; 16]. These techniques are grounded in contrastive learning [55; 60; 37], mask modeling [61], and other pretext tasks [3; 64]. PPKT [34] utilized the InfoNCE loss to facilitate the 3D network in distilling rich knowledge from the 2D image backbone. Sautier et al.[43] pioneered a superpixelto-superpoint contrastive loss for self-supervised 2D-to-3D representation distillation. Building on this, Mahmoud et al. [36] enhanced the approach by incorporating a semantically tolerant contrastive constraint and a class-balancing loss. Liu et al.[33] further refined these techniques through semantic-aware spatial and temporal consistency regularization, advancing feature learning. Moreover, Zhang et al. [62] explored cross-modality contrastive distillation across not only the image plane but also the bird-eye views.
Visual Foundation Models. The advent of powerful visual neural networks, trained on extensive datasets [39; 24] or through cutting-edge self-supervised learning techniques [6; 10; 19], catalyzed significant advancements within the community. Notably, the Segment Anything Model (SAM) [24] initiated a new paradigm in general-purpose image segmentation, demonstrating remarkable zero-shot transfer capabilities across a variety of downstream tasks. Building on this, Grounded-SAM [40] enhanced the model by incorporating elements from Grounding-DINO [32], an open-set object detector capable of recognizing and classifying previously unseen objects during training [32]. In our work, we leverage the inherent semantic awareness of these VFMs to generate weak semantic labels, which are crucial for our supervised contrastive distillation framework.
3D Scene Understanding. Traditional approaches to 3D scene understanding primarily utilize paradigms based on raw points [47; 14; 7; 53], voxels [15; 12], range views [25; 48], and multi-view fusion [17; 56]. Despite the effectiveness of these methods in capturing detailed environmental features, they are significantly constrained by their dependence on large-scale annotated data. Acquiring and annotating this data is both time-consuming and costly, limiting the scalability of 3D segmentation models [31]. To reduce the reliance on large annotated datasets, recent studies have also turned to semi-supervised [26; 20], weakly-supervised [30; 13], and active learning techniques [35; 54].
3 Proposed Method
Figure 2: The overall pipeline of our proposed OLIVINE. The pipeline starts with feature extraction via a trainable 3D backbone and a pre-trained 2D backbone, followed by feature alignment in a common space. The learning is driven by weakly-supervised contrastive distillation with coarse semantic labels, self-supervised distillation of randomly sampled point-pixel pairs, and semantic consistency regularization through the von Mises-Fisher distribution. Besides, our approach is also characterized by the novel sampling strategy of point-pixel pairs addressing spatial and category distribution imbalances.
Overview. As depicted in Fig. 2, our method, namely OLIVINE, integrates a visual foundation model for pre-training with paired point clouds and images. Feature extraction is conducted using a trainable 3D backbone for point clouds and a pre-trained 2D backbone for images, with these features then mapped into a common feature space via decoupled projection heads for both point-pixel level and category-level contrastive distillation. The representation learning in OLIVINE is driven by three objectives: weakly-supervised contrastive distillation using coarse semantic labels to identify positive pairs by category, self-supervised contrastive distillation applied to randomly sampled point-pixel pairs, and a regularization framework based on the von Mises-Fisher distribution to ensure semantic consistency. Additionally, we address imbalances in spatial distribution and category frequency through a targeted sampling strategy, ensuring a balanced representation in the learning process.
Notation. Let $P=\{p_{1},p_{2},...,p_{N}|p_{i}\in\mathbb{R}^{3}\}$ be a point cloud consisting of $N$ points collected by a LiDAR sensor, and =$\mathcal{I}=\{I_{c}\|\ c=1,...,N_{\mathrm{cam}}\}$ multi-view images captured by $N_{\mathrm{cam}}$ synchronized cameras, where $I_{c}\in\mathbb{R}^{H\times W\times 3}$ is a single image with height $H$ and width $W$.
3.1 Baseline Architecture
We follow the existing work [34] to perform the basic point-to-pixel contrastive distillation, based on which we build our whole pipeline. Starting with point cloud and image inputs, we employ distinct encoders for feature extraction. The 3D features are extracted using an encoder $f_{\mathrm{3D}}(\cdot):\mathbb{R}^{N\times 3}\to\mathbb{R}^{N\times C_{\mathrm{3D}}}$, which processes the point clouds to produce features of dimension $C$ per point. For image features, we use an encoder $f_{2D}(\cdot):\mathbb{R}^{H\times W\times 3}\to\mathbb{R}^{H^{\prime}\times W^{\prime}\times C_{\mathrm{2D}}}$, initialized with weights from pre-trained image models. This setup facilitates knowledge transfer from the 2D domain to the 3D domain through contrastive learning. For the computation of contrastive loss, we design trainable projection heads, $h_{\mathrm{2D}}^{\mathrm{pp}}$ for 2D features and $h_{\mathrm{3D}}^{\mathrm{pp}}$ for 3D features, both aligning the features into a unified dimensional space, $C$. Specifically, the 3D projection head $h_{\mathrm{3D}}^{\mathrm{pp}}$ is a linear layer with $\ell_{2}$-normalization, converting 3D features to a normalized $C$-dimensional space. Similarly, the 2D projection head $h_{\mathrm{2D}}^{\mathrm{pp}}$, a convolution layer with a 1$\times$1 kernel followed by a bi-linear interpolation layer adjusting the spatial dimensions by a factor of 4, also applies $\ell_{2}$-normalization.
Utilizing the calibration matrix, we establish dense point-to-pixel correspondences $\{\bm{F}^{\mathrm{3D}}_{i},\bm{F}^{\mathrm{2D}}_{i}\}_{i=1}^{M}$, where $\bm{F}^{\mathrm{3D}}_{i}$ and $\bm{F}^{\mathrm{2D}}_{i}$ represent the paired features of points and images for the $i$-th pair, with $M$ denoting the total count of such valid pairs. Previous methods achieve cross-modal knowledge transfer by attracting positive pairs and repelling negative pairs within the feature space, employing the InfoNCE loss [38]. The point-pixel level contrastive loss is defined as
$\mathcal{L}_{\mathrm{PPNCE}}=-\frac{1}{M_{s}}{\sum_{i=1}^{M_{s}}\log\left[\frac{\mathrm{exp}{(\langle\bm{F}^{\mathrm{3D}}_{i},\bm{F}^{\mathrm{2D}}_{i}\rangle/\tau)}}{\sum_{j=1}^{M_{s}}\mathrm{exp}{(\langle\bm{F}^{\mathrm{3D}}_{i},\bm{F}^{\mathrm{2D}}_{j}\rangle/\tau)}}\right]},$ (1)
[…]
3.2 Weakly-supervised Contrastive Distillation
Existing methods often mistakenly treat unmatched points and pixels that share semantic labels as negative pairs. This practice overlooks the intrinsic semantic connections within the same categories, leading to potential conflicts in the learning process where beneficial relationships are ignored. To address this, we utilize the Segment Anything Model, which adeptly interprets and translates semantic cues from text prompts into precise semantic segmentation of images. This application of SAM allows us to generate high-quality semantic labels without repetitive training, enhancing the learning process. We represent these labels as $Y^{\mathrm{co}}=\{y^{\mathrm{co}}_{i}\}_{i=1}^{M}$, where each label corresponds to a specific point-pixel pair.
In point-pixel level contrastive loss, the pixels that belong to the same category but do not correspond to the given anchor point are taken as negative samples (see Eq. (1)). Therefore, we argue that the 2D and 3D features used for weakly-supervised contrastive learning, which take the category information into consideration, should differ from the features $\bm{F}^{\mathrm{3D}}$ and $\bm{F}^{\mathrm{2D}}$ that represent the individual points and pixels. To address this issue, we apply another two heads $h_{\mathrm{2D}}^{\mathrm{sem}}:\mathbb{R}^{H^{\prime}\times W^{\prime}\times C_{\mathrm{2D}}}\to\mathbb{R}^{H\times W\times C}$ and $h_{\mathrm{3D}}^{\mathrm{sem}}:\mathbb{R}^{N\times C_{\mathrm{3D}}}\to\mathbb{R}^{N\times C}$ to extract the semantic-level feature embeddings $\bm{G}^{\mathrm{2D}}$ and $\bm{G}^{\mathrm{3D}}$.
For the sampled points and pixels, we use their semantic labels to identify positive and negative pairs. Positive pairs are defined as point and pixel features that share the same semantic label, whereas negative pairs are those with different labels [23]. The weakly-supervised contrastive loss is defined as
$\mathcal{L}_{\mathrm{sup}}=-\frac{1}{M_{s}}\sum_{i=1}^{M_{s}}\log\left[\frac{1}{|A(i)|}\sum_{a\in A(i)}\frac{\mathrm{exp}{(\langle\bm{G}^{\mathrm{3D}}_{i},\bm{G}^{\mathrm{2D}}_{a}\rangle/\tau)}}{\sum_{j=1}^{M_{s}}\mathrm{exp}{(\langle\bm{G}^{\mathrm{3D}}_{i},\bm{G}^{\mathrm{2D}}_{j}\rangle/\tau)}}\right],$ (2)
where $A(i)$ denotes the set of indices of matched point-to-pixel pairs in the batch that have the same class with $i$-th point-pixel pair, and $|A(i)|$ indicates its cardinality.
3.3 Semantic-guided Consistency Regularization
We advocate that the construction of latent semantic structures in feature space could enhance representation learning. By leveraging semantic labels derived from SAM inference, we organize points with identical semantic labels into coherent groups. This grouping promotes feature consistency within these semantic categories, thereby stabilizing the learning of feature representations across varied data instances and yielding structured feature space.
Distribution Assumption. Intuitively, the point features extracted by the projection head $h_{\mathrm{3D}}^{\mathrm{sem}}$ from the same class should exhibit similarity in the feature space. For the purposes of contrastive learning, these features are normalized to exist on the unit hypersphere. Consequently, we model the point features of each class $k$ as a von Mises-Fisher (vMF) distribution $\mathrm{vMF}(z;\mu_{k},\kappa_{k})$. This distribution is a spherical adaptation of the normal distribution, suitable for data constrained to a hypersphere [28]. Here, $\mu_{k}$ represents the mean direction, while $\kappa_{k}$ is the concentration parameter, indicating the degree to which category features are concentrated around $\mu_{k}$. The probability density function for the vMF distribution, applicable to a random $C$-dimensional unit vector $\bm{z}$, is formulated as follows:
$f_{C}(\bm{z};{\mu}_{k},\kappa_{k})=\mathcal{K}_{C}(\kappa_{k})\exp({\kappa_{k}\mu_{k}^{\top}\bm{z}}),$ (3)
where $\kappa\geq 0$ and $\|\mu\|_{2}=1$. The normalization constant $\mathcal{K}_{C}(\kappa)$ is defined as:
$\mathcal{K}_{C}(\kappa_{k})=\frac{\kappa_{k}^{C/2-1}}{(2\pi)^{C/2}\mathcal{I}_{(C/2-1)}(\kappa_{k})},$ (4)
$\mathcal{I}_{(C/2-1)}(x)=\sum_{m=0}^{\infty}\frac{1}{m!\Gamma(C/2-1+m+1)}\left(\frac{x}{2}\right)^{2m+C/2-1},$ (5)
where $\mathcal{I}_{(C/2-1)}$ is the modified Bessel function of the first kind at order $C/2-1$. The distribution exhibits a higher concentration around the mean direction $\mu_{k}$ as $\kappa_{k}$ increases, and becomes uniform on the hypersphere when $\kappa_{k}=0$.
Parameter Updating. Specifically, we conduct the semantic-guided consistency regularization in an two-stage framework. First, we update the parameter of $\mathrm{vMF}(z;\mu_{k},\kappa_{k}))$ with the features $\bm{G}^{\mathrm{3D}}$ extracted from point cloud branch. During training, we obtain the statistical value of feature embeddings via the EMA (Exponential Moving Average) algorithm, following:
$\bar{z}_{k}^{t}=\alpha\bar{z}_{k}^{t-1}+(1-\alpha)\bar{z}_{k}^{\prime t},$ (6)
where $\bar{z}_{k}^{\prime t}=\frac{1}{M_{k}}\sum_{i=1}^{M}\mathbf{1}_{\{y^{\mathrm{co}}_{i}=k\}}\bm{G}^{\mathrm{3D}}_{i}$ denotes the sample mean of class $k$ at $t$-th mini-batch, $\alpha$ is the fixed smoothness coefficient. The maximum likelihood estimates of the mean direction $\mu_{k}$ is simply the normalized arithmetic mean:
$\mu_{k}=\bar{z}_{k}/\bar{R}_{k},$ (7)
where $\bar{R}_{k}=||\bar{z}_{k}||$. And the concentration parameter $\kappa_{k}$ could be obtained by the solving the equation:
$A_{C}(\kappa)=\bar{R}_{k},$ (8)
where $A_{C}(\kappa)=\mathcal{I}_{C/2}(\kappa)/\mathcal{I}_{C/2-1}(\kappa)$. Sra [45] proposed a simple method to estimate the $\kappa_{k}$:
$\hat{\kappa}_{k}=\frac{\bar{R}_{k}(C-\bar{R}_{k}^{2})}{1-\bar{R}_{k}^{2}}.$ (9)
And we model the features of each observed point as a Spherical Dirac Delta distribution during training:
$\delta(z-\bm{G}^{\mathrm{3D}}_{i})=\left\{\begin{array}[]{l l}{0}&{z\not=\bm{G}^{\mathrm{3D}}_{i}}\\ {\infty}&{z=\bm{G}^{\mathrm{3D}}_{i}}\end{array}\right.$ (12)
Loss Function of Regularization. At the second step, we could perform the semantic-guided consistency regularization by pulling the points features and distribution of its corresponding category $\mathrm{vMF}(z;\mu_{k},\kappa_{k})$ with Kullback-Leibler (KL) Divergence loss:
$\mathcal{L}_{\mathrm{kl}}$ $=\frac{1}{M}\sum_{i=1}^{M}D_{\mathrm{KL}}(\delta(z-\bm{G}^{\mathrm{3D}}_{i})||\mathrm{vMF}(z;\mu_{k},\kappa_{k}))=\frac{1}{M}\sum_{i=1}^{M}\sum_{k=1}^{K}\mathbf{1}_{\{y^{\mathrm{co}}_{i}=k\}}-\log f_{C}(\bm{G}^{\mathrm{3D}}_{i};\mu_{k},\kappa_{k})$
$=\frac{1}{M}\sum_{i=1}^{M}\sum_{k=1}^{K}\mathbf{1}_{\{y^{\mathrm{co}}_{i}=k\}}-\log\mathcal{K}_{C}(\kappa_{k})-\kappa_{k}\mu_{k}^{\top}\bm{G}^{\mathrm{3D}}_{i}.$ (13)
In summary, the overall loss function for pre-training is written as $\mathcal{L}=\lambda_{1}\mathcal{L}_{\mathrm{PPNCE}}+\lambda_{2}\mathcal{L}_{\mathrm{sup}}+\lambda_{3}\mathcal{L}_{\mathrm{kl}}$, where $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are the weights to balance the three terms.
3.4 Density and Category-aware Sampling Strategy
Previous methods [34] neglect the spatial distribution and category frequency imbalances in the sampling of point-pixel pairs for contrastive distillation. To overcome these challenges, we introduce a novel sampling strategy that utilizes both the distance of each point from the LiDAR sensor and the occurrence frequency of its category. First, we calculate the distance of each point in the point cloud from the LiDAR sensor. We then apply kernel density estimation (KDE) to these distances to determine the probability density function of the spatial distribution of points. Given a point, its density can be calculated using the formula based on its distance $d$ from the LiDAR sensor:
$f_{\mathrm{h}}(d)=\frac{1}{Mh}\sum_{i=1}^{M}K_{h}\left(\frac{d-d_{i}}{h}\right),$ (14)
where $d_{i}$ denotes the distance of point $p_{i}$ from the LiDAR sensor, $h$ is the bandwidth, $K_{h}$ is the kernel function. This density estimation helps us understand how densely points are distributed with respect to their distance from the sensor, which is crucial for addressing areas of high point concentration that may bias the learning process.
Simultaneously, we assess the frequency of each category in the valid point-pixel pairs. By counting the occurrences of each category, we can identify which categories are underrepresented or overrepresented in the dataset.
Combining these two dimensions of analysis, we define the sampling probability for each point as inversely proportional to both its KDE-derived density and its category occurrence frequency. Mathematically, the sampling probability for a point $p_{i}$ is given by:
$\rho(p_{i})=1/(f_{\mathrm{h}}(d_{i})\times|A(i)|).$ (15)
By implementing this sampling strategy, we aim to ensure a more uniform representation of both spatial and categorical dimensions in our contrastive learning setup. This method reduces the risk of overfitting to dense clusters of points or to frequently occurring categories, thereby fostering a more robust and generalizable representation learning.
4 Experiments
4.1 Experimental Setup of 3D Pretraining
Network Architectures. For the image processing branch, we utilize the ResNet-50 structure as the core architecture. This 2D backbone is initialized with weights that have been pre-trained using MoCov2 [9] on the ImageNet dataset. To preserve the receptive field while keeping the spatial resolution intact, we substitute the second and subsequent stridden convolutions with dilated convolutions, following established methodologies [43]. The up-sampling projection head includes a 1$\times$1 convolutional layer that reduces the channel count from 2048 to 64, followed by a bi-linear interpolation up-sampling layer that enlarges the scale by a factor of 4. This up-sampling process effectively restores the resolution of the 2D feature map to match that of the original input images, specifically to the size of 416$\times$224.
In the point cloud processing branch, we adopt two types of backbones. For the 3D semantic segmentation task, we employ the Sparse Residual 3D U-Net 34 (SR-UNet34) [41], adhering to practices previously established in SLidR [43]. The output from SR-UNet34 offers 256 channels, whereas the image branch outputs a 64-dimensional feature map. To align these dimensions, a 3D convolutional layer is used in the projection head to reduce the channel count of the point features to 64. We process the 3D point data into voxels to serve as input for the SR-UNet. The voxels are formatted in Cartesian coordinates covering an X-axis and Y-axis range of [-51.2m, 51.2m] and a Z-axis range of [-5.0m, 3.0m], with each voxel measuring (0.1m, 0.1m, 0.1m). To fully evaluate our method, we pre-train and transfer another VoxelNet [70] for the 3D object detection task.
Pre-training Details. We utilize momentum SGD for optimization, setting the initial learning rate at 0.5 and 0.01 for SR-UNet34 and VoxelNet respectively, with a momentum of 0.9 and a weight decay of 1e-4. To adjust the learning rate, we employ a cosine annealing scheduler [4] that gradually reduces it from the initial value to 0 over 50 epochs. The 3D network is pre-trained for these 50 epochs on four NVIDIA-3090 GPUs, processing a total batch size of 16, unless specified otherwise. For data augmentation, we incorporate several techniques. For the point cloud data, we apply random rotations around the z-axis, randomly flip the x and y axes, and omit points within a randomly selected cuboid, following the method described in [68]. For image data, augmentations include random horizontal flips and random crop-resize operations. In terms of generating weak semantic labels, the prompts we provided to the Grounded-SAM [40] encompass a total of 13 object categories: barrier, bicycle, bus, car, motorcycle, pedestrian, traffic cone, truck, road, sidewalk, terrain, vegetation, and building.
Table 1: Comparison of various pre-training techniques for semantic segmentation tasks using either finetuning or linear probing. This evaluation uses different proportions of accessible annotations from the nuScenes datasets and presents the mean Intersection over Union (mIoU) scores on the nuScenes validation set.
  Initialization Present at Lin. prob. Finutune
100% 1% 5% 10% 25% 100%
Random - 8.1 30.30 47.84 56.15 65.48 74.66
PointContrast [55] ECCV’20 21.90 32.50 - - - -
DepthContrast [68] ICCV’21 22.10 31.70 - - - -
PPKT [34] Arxiv’21 35.90 37.80 53.74 60.25 67.14 74.52
SLidR [43] CVPR’22 38.80 38.30 52.49 59.84 66.91 74.79
ST-SLidR [36] CVPR’23 40.48 40.75 54.69 60.75 67.70 75.14
HVDistill [62] IJCV’24 39.50 42.70 56.60 62.90 69.30 76.60
Seal [33] NeurIPS’23 44.95 45.84 55.64 62.97 68.41 75.60
Ours - 47.30 46.12 57.51 63.04 69.39 76.13
4.2 Transfer on Semantic Segmentation
Evaluation Protocol. We evaluate the learned representations for semantic segmentation on nuScenes-lidarseg and SemanticKITTI datasets. The nuScenes-lidarseg and SemanticKITTI datasets contain 16 and 19 semantic categories for validation, respectively. We modify the network by adding a 3D convolutional layer to the pre-trained backbone as the segmentation head. Basically, we finetune the whole network on different portions of annotated data. Following previous works [43; 36], we finetune the network for 100 epochs with a batch size of 10 and 16 on SemanticKITTI and nuScenes-lidarseg, respectively. The initial learning rate of the backbone and the segmentation head are set to 0.05 and 2.0, respectively. When utilizing 1% of the annotated data, the network is fine-tuned for 100 epochs, whereas for other percentages, it is fine-tuned for 50 epochs. In another protocol, we evaluate the quality of learned representation by linear probing. Different from the setting of finetuning, we optimize only the added segmentation head and keep the weights of backbone $f_{\mathrm{3D}}$ frozen on the nuScenes-lidarseg dataset. For both protocols, the training objective is a linear combination of the cross-entropy loss and the Lovász-Softmax loss [2].
Figure 3: The visual results of various point cloud pretraining strategies, pre-trained on nuScenes and fine-tuned using merely 1% of annotated data, are displayed. To illustrate the distinctions, we mark correctly predicted areas in gray color and incorrect ones in red.
Table 2: Comparison of various pre-training techniques for semantic segmentation through finetuning on the SemanticKITTI dataset with only 1% annotated data.
  Initialization Random Point- Contrast Depth- Contrast PPKT SLidR ST-SLidR Seal Ours
mIoU (%) 39.50 41.10 41.50 44.00 44.60 44.72 46.63 47.35
Improvement - +0.60 +1.00 +4.50 +5.10 +5.22 +7.13 +7.85
Table 3: Finetuning results on SemanticKITTI across various percentages of annotated data. The table compares the improvement achieved by our method relative to the SLidR [43].
  Initialization 1% 5% 10% 20% 100%
Random 39.5 - 45.7 - 51.5 - 56.0 - 56.9 -
SLidR [43] 44.6 +5.1 54.7 +9.0 56.3 +4.8 56.7 +0.7 57.1 +0.2
Ours 47.4 +7.9 56.5 +10.8 58.2 +6.7 59.5 +3.5 59.7 +2.8
Results of Linear Probing. Under the linear probing scenario, our method achieves the highest mIoU of 47.30%, surpassing the previous state-of-the-art method, Seal [33], which records a mIoU of 44.95% (see Table 1). This performance indicates a significant improvement in extracting useful features directly from pre-trained models without additional training of the 3D backbone.
Results of Fine-tuning. For the fine-tuning on nuScenes, our consistently excel, particularly at smaller data proportions. With only 1% of the training data of the nuScenes, our method achieves a mIoU of 46.12%, outperforming the HVDistill [62] that had a mIoU of 42.70%. This trend persists across other data proportions, with our method consistently leading or closely competing with the best results, particularly at 57.51% mIoU with 5% of the data and 63.04% mIoU with 10% of the data. The qualitative results are presented in Fig. 3.
As shown in Table 2, our method also achieves state-of-the-art performance on the SemanticKITTI dataset. With only 1% of the data annotated, our method achieved a mIoU of 47.35%, which is a substantial improvement of 7.85% over the baseline. We also fine-tuned various models on the SemanticKITTI dataset to assess their performance across a spectrum of annotated data availability, from as low as 1% to full data utilization (see Table 3). Specifically, with only 1% of the data annotated, our method achieved a mIoU of 47.35%, which is a substantial improvement of 7.85% over the baseline. This trend continues as more data becomes available; with 5% of data, our model further increased its lead by achieving a mIoU of 56.5%, reflecting an improvement of 10.8%.
Table 4: Comparison of our method with other pre-training techniques through fine-tuning on the KITTI dataset. The results reflect the 3D object detection performance under moderate difficulty on the validation set.
  Detectors Initialization Car Pedestrian Cyclist mAP@40
SECOND [59] Random 81.5 50.9 66.5 66.3
PPKT [34] 81.8 51.4 68.2 67.1
SLidR [43] 81.9 51.6 68.5 67.3
Ours 82.0 53.2 69.8 68.3
PV-RCNN [44] Random 84.5 57.9 71.3 71.3
STRL [22] 84.7 57.8 71.9 71.5
PPKT [34] 83.2 55.5 73.8 70.8
SLidR [43] 84.4 57.3 74.2 71.9
Ours 84.8 59.3 74.2 72.8
4.3 Transfer on 3D Object Detection
Evaluation Protocol. In evaluating our pre-trained model for 3D object detection, we utilized two prominent architectures: SECOND [59] and PV-RCNN [44]. Both are built on the VoxelNet 3D backbone [70], which processes voxels via 3D sparse convolutions and includes a 2D backbone for bird’s-eye-view encoding after BEV projection. The primary distinction between the architectures is in their detection heads. SECOND uses a region proposal network (RPN) directly on the 2D backbone, while PV-RCNN refines RPN predictions with fine-grained keypoint features, enhancing bounding box accuracy and confidence in estimations.
In the fine-tuning stage, we integrate the detection head of SECOND or PV-RCNN with the pre-trained backbone (VoxelNet). This integrated detector is then fine-tuned on the train data of KITTI [18], which includes implementations of these detectors and follows the standard training parameters specified by OpenPCDet [46]. We conduct the fine-tuning three separate times, and report the highest mean Average Precision (mAP) recorded on the KITTI validation set.
Results. The experimental results detailed in Table 4 showcase the performance of various initialization strategies. When using the SECOND architecture, our method outperforms other pre-training techniques. Starting from a baseline with random initialization, the performance improves consistently with more specialized pre-trained weights like PPKT and SLidR, and ultimately, our method achieves the highest mAP at 68.3%. Significant gains are observed across all categories, with particularly notable improvements in detecting pedestrians and cyclists. Similarly, the PV-RCNN architecture benefits from refined initialization methods. Our method again yields the highest overall mAP@40 at 72.8%, surpassing the performance of SLidR.
Table 5: Ablation study of each component pre-trained on nuScenes and fine-tuned on nuScenes-lidarseg and SemanticKITTI. WCD: Weakly-supervised Contrastive Distillation. SCR: Sematic-guided Consistency Regularization. DCAS: Denstiy and Category-aware Sampling.
  Exp. WCD SCR DCAS nuScenes SemanticKITTI
LP 1% 5% 10% 25% 100% 1%
1 $\times$ $\times$ $\times$ 36.87 37.89 53.15 60.33 67.03 74.59 44.12
2 $\times$ $\times$ 42.10 42.33 55.22 61.53 67.70 75.06 45.58
3 $\times$ $\times$ 40.58 41.99 54.49 60.98 67.69 74.88 45.67
4 $\times$ 44.76 44.91 56.01 63.12 68.74 75.48 46.60
5 $\times$ 44.15 43.96 55.75 62.49 68.13 75.07 46.07
6 47.30 46.12 57.51 63.04 69.39 76.13 47.35
Table 6: Comprehensive ablation studies for the key components. We report the fine-tuned results on nuScenes-lidarseg and SemanticKITTI (S.K.) datasets with 1% of the labeled data.
(a) Different types of supervision utilized for contrastive distillation.
  Method nuScenes S.K.
Baseline 37.89 44.12
w/ Weak label 46.12 47.35
w/ GT label 57.72 51.56
(b) Different architectures of projection heads.
  Heads nuScenes S.K.
Not Decoupled 42.59 45.46
Decoupled 46.12 47.35
Improvement +3.53 +1.89
(c) Different distributions to model semantic features.
  Distribution nuScenes S.K.
Deterministic 44.15 45.98
vMF 46.12 47.35
Improvement +1.97 +1.37
(d) Different sampling strategies.
  Sampling nuScenes S.K.
Random 44.91 46.60
Density-aware 45.33 46.77
Category-aware 45.74 46.82
DCAS 46.12 47.35
4.4 Ablation Studies
Effect of Key Components. In Table 5, we investigate the effect of each added component in our method. The integration of Weakly-supervised Contrastive Distillation alone yields a significant increase in performance, improving mIoU by 5.23% for linear probing. Similarly, incorporating Sematic-guided Consistency Regularization also enhances model performance, delivering a 3.71% increase in mIoU. When these components are combined, they synergistically contribute to a further mIoU gain of 7.89% for linear probing. Additionally, the application of Denstiy and Category-aware Sampling independently provides a substantial performance boost. The culmination of integrating all proposed components results in the optimal model, achieving a mIoU improvement of 10.43% for linear probing. This comprehensive analysis underscores the effectiveness of each component and their collective impact in enhancing the model’s segmentation capabilities.
Potential of Supervised Contrastive Distillation. As mentioned in Section. 3, we perform weakly-supervised contrastive distillation with the pseudo-labels predicted by SAM. With the free and available models, our method learns effective 3D representations. When we replace the weak labels with the ground truth provided in nuScenes-lidarseg dataset, we can obtain a significant improvement in downstream tasks (see Table 6(a)). The results further demonstrate the effectiveness of supervision for cross-modal contrastive distillation and the potential of our pipeline with stronger VFMs.
Effect of the Decoupled Projection Heads. The experimental results outlined in Table 6(b) demonstrate the effectiveness of employing decoupled projection heads in our model. These results highlight a distinct performance enhancement when projection heads are specialized for distinct tasks — specifically, self-supervised and weakly-supervised contrastive distillation. On the nuScenes dataset, the implementation of decoupled projection heads results in a mIoU improvement of 3.53%, indicating a robust enhancement in the model’s ability to generalize from the training data. Similarly, for the SemanticKITTI dataset, a gain of 1.89% in mIoU is observed, further substantiating the benefits of this architectural modification.
Effect of the vMF Distribution. The ablation study in Table 6(c) compares the use of deterministic (Dirac delta) and von Mises-Fisher (vMF) distributions for modeling semantic features of each class, demonstrating clear advantages with vMF on both nuScenes and SemanticKITTI datasets. Specifically, the vMF distribution, with its adjustable concentration parameter, provides a mIoU improvement of 1.97% on nuScenes and 1.37% on SemanticKITTI compared to the deterministic approach. The learned concentration parameter that represents the uncertainty helps in mitigating overfitting by providing robustness against inaccuracies in coarse semantic labels.
4.5 Visualization and Analysis
The T-SNE visualization shown in Fig. 4 demonstrates the efficacy of our method in achieving a more discriminative and well-separated feature distribution for each category compared to the baseline model PPKT [34]. To some extent, each category forms a distinct cluster with our method, with relatively clear boundaries separating different classes. This enhanced clustering effect underscores the benefits of our approach, which incorporates semantic supervision and applies semantic-guided consistency regularization during pre-training.
Figure 4: T-SNE visualization of the extracted point cloud features by PPKT and our OLIVINE (with head $h_{\mathrm{3D}}^{\mathrm{sem}}$). Each feature is colorized based on its ground-truth semantic labels on nuScenes dataset.
5 Conclusion
We have introduced OLIVINE, a novel method that leverages visual foundation models for fine-grained image-to-LiDAR knowledge transfer. The key ingredient of our method is utilizing the weak semantic labels generated from VFMs to avoid semantically similar negative samples and tackle the challenge of “self-conflict” challenges. We further exploit the semantic labels with semantic-guided consistency regularization, which makes embeddings of points in the same class remain consistent across varying inputs and yields structured feature space. Extensive experiments on various datasets confirm that our method achieves superior performance in downstream tasks compared to existing image-to-LiDAR contrastive distillation methods.
Appendix A Appendix
In this appendix, we present the details omitted from the manuscript due to space constraints. The appendix is organized as follows:
  • Section A.1: Details of the dataset and evaluation metrics.
  • Section A.2: More quantitative results.
  • Section A.3: More qualitative results.
  • Section A.4: Potential limitations of our method.
  • Section A.5: Societal and environmental impact of our work.
  • Section A.6: Public resources used in this work.
A.1 Dataset and Evaluation Metric
NuScenes Dataset. The NuScenes dataset, compiled from driving recordings in Boston and Singapore, utilizes a vehicle equipped with a 32-beam LiDAR and additional sensing technologies [5]. This comprehensive dataset is equipped with the typical sensor array found on autonomous vehicles, including a 32-beam LiDAR setup, six cameras, and radar systems, ensuring full 360-degree environmental perception. It includes 850 driving scene snippets, with 700 designated for training and 150 for validation, each scene lasting 20 seconds with annotations provided every 0.5 seconds. The dataset features extensive annotations across several object categories, including vehicles, pedestrians, bicycles, and road barriers, with each object encapsulated in a 3D bounding box and supplemented with attributes detailing visibility, activity, and pose.
NuScenes-lidarseg Dataset. The nuScenes dataset now encompasses features for semantic and panoptic segmentation through its extension, nuScenes-lidarseg [5]. This enhanced dataset provides semantic labeling across 32 distinct categories, with each point in the dataset’s keyframes meticulously annotated. We utilize the 700 training scenes equipped with segmentation labels for refining our semantic segmentation models, and we assess model performance using the 150 scenes in the validation set.
SemanticKITTI Dataset. The SemanticKITTI (SK) dataset features paired RGB images and point cloud data derived from KITTI’s urban scenes, specifically designed for semantic segmentation tasks [1]. This dataset is gathered using sensors mounted on a vehicle, including more than 200,000 images alongside their corresponding point clouds across 21 distinct sequences. Both images and point clouds are aligned to maintain a consistent relative transformation. Originally, the images are captured at a resolution of 1241x376 pixels, and each point cloud is composed of roughly 40,000 3D points. In line with standard practices, the dataset divides into training and validation sets, with 10 sequences designated for training and the eighth sequence reserved for validation.
KITTI Dataset. KITTI is a crucial dataset for advancing 3D object detection in autonomous driving. With 7481 training and 7518 test point clouds, it covers diverse urban and suburban environments [18]. The dataset includes 3D point clouds and RGB images captured using a Velodyne HDL-64E LiDAR sensor. Calibration information between the camera and LiDAR is provided, essential for cross-modal knowledge transfer or sensor fusion tasks. Annotated with 3D bounding boxes, it features common objects like cars, pedestrians, and cyclists. The dataset is split into training (3712 samples) and validation (3769 samples) subsets.
Evaluation Metrics. In semantic segmentation tasks, performance is assessed through Intersection-over-Union (IoU) for individual classes and mean IoU (mIoU) across all classes. In 3D object detection, the 3D detector’s efficacy on the KITTI dataset is measured using Average Precision (AP) metrics at IoU thresholds of 0.7 for cars, 0.5 for pedestrians, and 0.5 for cyclists. Similarly, for the Waymo dataset, evaluation is based on 3D mean Average Precision (mAP).
Figure 5: Class distribution at the pixel level for nuScenes dataset.
Figure 6: Class distribution at the pixel level for SemanticKITTI dataset.
A.2 More Quantitative Results
Class Unbalance in Datasets. The visualizations in Figures 5 and 6 illustrate the class distribution at the pixel level for the nuScenes and SemanticKITTI datasets, respectively. These figures reveal a significant class imbalance, a common challenge in many real-world datasets, where some classes are overwhelmingly more frequent than others. Such imbalance can skew the training process, leading to models that perform well on frequent classes but poorly on rare ones. This disparity predominantly affects the model’s ability to generalize effectively across different scenarios, particularly underrepresented ones, resulting in biased predictions and reduced overall accuracy. For instance, infrequent but critical objects like pedestrians or bicycles might not be detected reliably, which is particularly concerning in autonomous driving contexts where safety is paramount.
To mitigate these issues, our method incorporates an optimized sampling strategy. This strategy involves adjusting the probability of selecting samples from underrepresented classes during the training process. By increasing the likelihood of including rare classes in the training set, we ensure that the model does not overlook these important but less frequent categories.
Table 7: Per-class results on the nuScenes-lidarseg dataset using only 1% of the labeled data for fine-tuning. This chart displays the IoU scores for each category, with the highest and second-highest scores marked in dark blue and light blue, respectively.
  Method barrier bicycle    bus    car const. veh. motor pedestrian traffic cone trailer truck driv. surf. other flat sidewalk terrain manmade vegetation mIoU
Random 0.0 0.0 8.1 65.0 0.1 6.6 21.0 9.0 9.3 25.8 89.5 14.8 41.7 48.7 72.4 73.3 30.3
PointContrast 0.0 1.0 5.6 67.4 0.0 3.3 31.6 5.6 12.1 30.8 91.7 21.9 48.4 50.8 75.0 74.6 32.5
DepthContrast 0.0 0.6 6.5 64.7 0.2 5.1 29.0 9.5 12.1 29.9 90.3 17.8 44.4 49.5 73.5 74.0 31.7
PPKT 0.0 2.2 20.7 75.4 1.2 13.2 45.6 8.5 17.5 38.4 92.5 19.2 52.3 56.8 80.1 80.9 37.8
SLidR 0.0 1.8 15.4 73.1 1.9 19.9 47.2 17.1 14.5 34.5 92.0 27.1 53.6 61.0 79.8 82.3 38.3
ST-SLidR 0.0 2.7 16.0 74.5 3.2 25.4 50.9 20.0 17.7 40.2 92.0 30.7 54.2 61.1 80.5 82.9 40.8
Ours 0.0 6.7 55.6 82.7 10.3 26.8 57.8 23.1 18.8 47.3 93.9 32.4 55.8 62.4 82.0 82.3 46.1
Table 8: Per-class performance on SemanticKITTI with 1% of the labeled data utilized for fine-tuning. The figure shows the IoU for each class, where the top and second top scores are indicated by dark blue and light blue backgrounds, respectively.
  Method    car bicycle motorcycle truck other-vehicle person bicyclist motorcyclist    road parking sidewalk other-ground building fence vegetation trunk terrain    pole traffic-sign mIoU
Random 91.2 0.0 9.4 8.0 10.7 21.2 0.0 0.0 89.4 21.4 73.0 1.1 85.3 41.1 84.9 50.1 71.4 55.4 37.6 39.5
PPKT 91.3 1.9 11.2 23.1 12.1 27.4 37.3 0.0 91.3 27.0 74.6 0.3 86.5 38.2 85.3 58.2 71.6 57.7 40.1 43.9
SLidR 92.2 3.0 17.0 22.4 14.3 36.0 22.1 0.0 91.3 30.0 74.7 0.2 87.7 41.2 85.0 58.5 70.4 58.3 42.4 44.6
Ours 92.7 12.3 19.0 30.2 18.3 45.6 31.4 0.1 91.8 32.1 75.3 0.6 87.8 45.0 84.9 58.7 71.4 59.6 42.8 47.4
Per-class Performance. In Tables 7 and 8, we showcase the per-class performance of various point cloud pretraining strategies, including our method and other baselines, fine-tuned using just 1% of labeled data from the nuScenes-lidarseg and SemanticKITTI datasets. Our approach consistently surpasses other methods, achieving the highest mean Intersection over Union (mIoU) in nearly all categories. This marked superiority is also significant in complex categories like the bus and truck that demand more precise segmentation, highlighting our method’s robustness in processing sparse and intricate data scenarios.
A.3 More Qualitative Results
Figure 7: Visualization of the similarities between image and point cloud feature. In the first column, we show the raw image and the projection of anchor point in the image. In second columns, we illustrate the similarities between 3D query and 2D features extracted by the conventional projection heads $h_{\mathrm{3D}}^{\mathrm{pp}}$ and $h_{\mathrm{2D}}^{\mathrm{pp}}$ for point-pixel level contrastive distillation. In third columns, we illustrate the similarities between 3D query and 2D features extracted by the extra projection heads $h_{\mathrm{3D}}^{\mathrm{sem}}$ and $h_{\mathrm{2D}}^{\mathrm{sem}}$ for weakly-supervised (category-aware) contrastive distillation.
2D-3D Feature Similarities. The visualization presented in Figure 7 showcases the feature similarities between image and point cloud data as extracted through different projection heads. In the first column, the raw image is displayed alongside the projection of an anchor point within that image, setting the context for the comparison. The second column visualizes the feature similarities as extracted by the conventional projection heads, $h_{\mathrm{3D}}^{\mathrm{pp}}$ and $h_{\mathrm{2D}}^{\mathrm{pp}}$, used for point-pixel level contrastive distillation. Here, only the features of the pixel that directly correspond to the anchor point show a high degree of similarity, emphasizing a tight, point-to-point correspondence.
In contrast, the third column introduces results from the extra projection heads, $h_{\mathrm{3D}}^{\mathrm{sem}}$ and $h_{\mathrm{2D}}^{\mathrm{sem}}$, which are designed for weakly-supervised (category-aware) contrastive distillation. This setup reveals a broader similarity pattern, where points and pixels sharing the same category exhibit notably higher feature similarities. This suggests that our newly proposed projection heads are effective in capturing and reinforcing category-level feature associations across the 3D and 2D domains, thus enhancing the model’s ability to recognize semantically similar but spatially disparate features.
Visual Results on Downstream Tasks. In Figures 8, 9, 10, and 11, we present additional qualitative results from fine-tuning tasks on downstream datasets. The application of pre-training strategies markedly improves model accuracy over baselines that use random initialization. Notably, our proposed OLIVINE outperforms SLiDR [43], highlighting its superior segmentation capabilities. Despite these advancements, we note the occurrence of false positive predictions in edge cases, which we aim to address in future research.
Visualization of the Weak Semantic Labels. The weak labels generated by Grounded-SAM [40] using targeted prompts play a vital role in our processing pipeline. While reviewing these labels, we observe instances of imprecision (see Fig. 12 and Fig. 13). Should future advancements in Segmentation Anything Models yield more robust and accurate results, the effectiveness of our 3D pre-training strategy is likely to improve significantly. This progression would enhance our model’s ability to interpret and learn from nuanced environmental data, ultimately leading to superior representation learning.
A.4 Potential Limitations
While our method, OLIVINE, effectively enhances the fine-grained image-to-LiDAR contrastive distillation process and demonstrates significant improvements in 3D scene understanding, there are technical and potential limitations that merit attention:
Semantic Label Accuracy. The accuracy of the weak semantic labels generated by VFMs is critical to the success of our model. Any deficiencies in these labels could propagate through the learning process, potentially compounding errors in the learned representations.
Training Data Diversity. Currently, our model is pre-trained using a single dataset, which may limit its applicability to environments or scenarios not well-represented in the training data. Expanding the training to include diverse datasets with varying characteristics could enhance the robustness and generalizability of our model.
Dependency on High-Quality Data Calibration. Our framework relies on the precise calibration and synchronization between LiDAR sensors and cameras. In real-world applications, perfect synchronization and calibration can be challenging to maintain, potentially affecting the accuracy and reliability of the semantic labels generated and subsequently the distillation process [43; 33; 62].
These limitations highlight areas for future development and research, suggesting a path towards more robust, adaptable, and efficient systems for 3D scene understanding in diverse and dynamic environments.
A.5 Societal and Environmental Impact
Our method, OLIVINE, which enhances image-to-LiDAR contrastive distillation using Visual Foundation Models, significantly impacts society and the environment. Societally, it boosts the safety and reliability of autonomous systems, increasing public trust and improving data analysis in various industries. Environmentally, deploying advanced deep learning models requires increased computational resource usage, which can lead to higher energy consumption and associated carbon emissions. This is particularly relevant during the intensive training phases that require high-performance GPUs and long training durations, especially as the model scales to larger datasets or more complex scenarios. Conversely, by distributing our pre-trained models, we aim to reduce the need for repetitive training across multiple downstream tasks, which can decrease the overall computational load and energy consumption needed to achieve high performance in various applications. This aspect potentially mitigates some of the environmental costs.
A.6 Public Resources Used
We acknowledge the use of the following public resources, during the course of this work:
Figure 8: Qualitative results of fine-tuning on 1% of the nuScenes-lidarseg dataset with different pre-training strategies. Note that the results are shown as error maps on the right, where red points indicate incorrect predictions. Best viewed in color and zoom in for more details.
Figure 9: Qualitative results of fine-tuning on 1% of the nuScenes-lidarseg dataset with different pre-training strategies. Note that the results are shown as error maps on the right, where red points indicate incorrect predictions. Best viewed in color and zoom in for more details.
Figure 10: Qualitative results of fine-tuning on 1% of the SemanticKITTI dataset with different pre-training strategies. Note that the results are shown as error maps on the right, where red points indicate incorrect predictions. Best viewed in color and zoom in for more details.
Figure 11: Qualitative results of fine-tuning on 1% of the SemanticKITTI dataset with different pre-training strategies. Note that the results are shown as error maps on the right, where red points indicate incorrect predictions. Best viewed in color and zoom in for more details.
Figure 12: Illustration of the weak semantic labels predicted by Grounded SAM. The top half of the figure displays the raw RGB images and LiDAR point clouds, while the bottom half presents the corresponding weak semantic labels applied to both images and point clouds, aligned using camera parameters. Each distinct segment is represented by a unique color. Best viewed in color.
Figure 13: Illustration of the weak semantic labels predicted by Grounded-SAM. The top half of the figure displays the raw RGB images and LiDAR point clouds, while the bottom half presents the corresponding weak semantic labels applied to both images and point clouds, aligned using camera parameters. Each distinct segment is represented by a unique color. Best viewed in color.